y''+4y=cos^2(x)


1
Expert's answer
2021-04-15T07:26:11-0400

Given: "y{}''+4y=cos^2x"

Require to solve the given differential equation.

Auxiliary equation corresponding to the given homogeneous differential equation "y{}''+4y=0"

is "m^2+4=0"


Solve the equation "m^2+4=0"

Now "m^2+4=0\Rightarrow m=\pm 2i"


General solution to the homogeneous differential equation is

"y_h=c_1cos2x+c_2sin2x"


Now let us find the particular solution.

Use the method of variation of parameters to find the particular solution.


Using the method of variation of parameters, Particular solution is "y_p=Au+Bv,=cos2x,v=sin2x" and

"A=-\int \frac{vg(x)}{uv{}'-vu{}'}dx,B=\int \frac{ug(x)}{uv{}'-vu{}'}dx" , where "g(x)=cos^2x"


First let us find "uv{}'-vu{}'"


Now "u=cos2x,v=sin2x\Rightarrow uv{}'-vu{}'=(cos2x)(2cos2x)-(sin2x)(-2sin2x)=2cos^2{2x}+2sin^2{2x}=2(1)=2"

Using the above, we get

"A=-\int \frac{vg(x)}{uv{}'-vu{}'}dx\Rightarrow A=-\int \frac{(sin2x)(cos^2x)}{2}dx=-\int \frac{(2sinxcosx)(cos^2x)}{2}dx"

"\Rightarrow A=-\int t^3(-dt),t=cosx\Rightarrow dt=-sinxdx"

Then, we get "A=-\frac{t^4}{4}=-\frac{cos^4x}{4}"


And

"B=\int \frac{ug(x)}{uv{}'-vu{}'}dx\Rightarrow B=\int \frac{(cos2x)(cos^2x)}{2}dx"

"\Rightarrow B=\frac{1}{2}\int(cos2x)\left [ \frac{1+cos2x}{2} \right ]dx"

"\Rightarrow B=\frac{1}{4}\int(cos2x)dx+\frac{1}{4}\int cos^2{2x}dx"

"\Rightarrow B=\frac{1}{4}\int(cos2x)dx+\frac{1}{4}\int \frac{1+cos4x}{2}dx"

"\Rightarrow B=\frac{1}{4}\left [ \frac{sin2x}{2} \right ]+\frac{1}{8}\left [ x+\frac{sin4x}{4} \right ]"

"\Rightarrow B=\frac{1}{8}sin2x+\frac{x}{8}+\frac{sin4x}{32}"


Using the above, we have

"y_p=-\frac{1}{4}(cos^4x)(cos2x)+\left [ \frac{1}{8}sin2x+\frac{x}{8}+\frac{sin4x}{32} \right ](sin2x)"


Therefore general solution is

"y=y_h+y_p"


"y=c_1cos2x+c_2sin2x-\frac{1}{4}(cos^4x)(cos2x)+\left [ \frac{1}{8}sin2x+\frac{x}{8}+\frac{sin4x}{32} \right ](sin2x)"